40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 18 23:59:53 2017, gautam, Update, ALS, X Arm ALS lock 
    Reply  Sat Aug 19 01:35:08 2017, ericq, Update, ALS, X Arm ALS lock ALS_comparison.pdf
    Reply  Mon Aug 21 23:38:55 2017, gautam, Update, ALS, ALS out-of-loop noise ALS_21082017.pdf
       Reply  Tue Aug 22 02:19:11 2017, gautam, Update, ALS, ALS OLTFs ALS_OLTFs.pdfALS_OLTF_Aug2017.zip
          Reply  Tue Aug 22 23:27:14 2017, rana, Update, ALS, ALS OLTFs 
Message ID: 13230     Entry time: Sat Aug 19 01:35:08 2017     In reply to: 13229
Author: ericq 
Type: Update 
Category: ALS 
Subject: X Arm ALS lock 

My motivation tonight was to get an up-to-date spectrum of a calibrated measurement of the out-of-loop displacement of an arm locked on ALS (using the PDH signal as the out-of-loop sensor) to compare the performance of ALS control noise with the Izumi et al green locking paper. 

I was able to fish out the PSD from the paper from the 40m svn, but the comparison as plotted looks kind of fishy. I don't see why the noise from 10-60Hz should be so different/worse. We updated the POX counts to meters conversion by looking at the Hz-calibrated ALSX signal and a ~800Hz line injected on ETMX.

Attachment 1: ALS_comparison.pdf  177 kB  | Hide | Hide all
ALS_comparison.pdf
ELOG V3.1.3-